Which of the following statements are true? Give reasons for your answers, in the form of a short proof or a counter example.
(1) d^2/dx^2=(dy/dx) ^2
(2) The inverse function of y=e^3x is y=1/3lnx.
(3) If f is increasing and f(x) >0 on an interval I, then g(x) = 1/f(x) is deceasing on I.
(4) An equation of the tangent line to the parabola y=x^2 at (-2, 4) is you=2x(x+2).
(5) If f is one - one onto and differentiable on R, then (f^-1) '(6) =1/f'(6) .
1
Expert's answer
2020-03-16T13:41:01-0400

(1) It's false. For example, if "y=x" then "\frac{d^2y}{dx^2}=0" and "(\frac{dy}{dx})^2=1"

(2) It's false. "\ln y=3x" hence "x=\frac{\ln y}{3}" then the inverse function "y=\frac{\ln x}{3}"

(3) It's true. If "f" is increasing hence "f'>0" then "g'(x)=(\frac{1}{f(x)})'=-\frac{f'(x)}{(f(x))^2}<0" hence "g(x)" is decresing on I

(4) It's false. The tangent line is "y-4=y'(x+2)" as "y'=2x" hence "y=2x(x+2)+4"

(5) It's false. For example "f=x" then "f^{-1}=\frac{1}{x}" then "f'(6)=1" and "(f^{-1})'(6)=-\frac{1}{36}"